# MonoRollBot: 3-DOF Spherical Robot with Underactuated Single Compliant Actuator Design

### v2

@@ Abstract
Spherical rolling robots have garnered significant attention in the field of mobile robotics for applications such as inspection and space exploration. Designing underactuated rolling robots poses challenges in achieving multi-directional propulsion with high degrees of freedom while utilizing a limited number of actuators. This paper presents the MonoRollBot, a novel 3-degree-of-freedom (DOF) spherical robot that utilizes an underactuated mechanism driven by only a single spring-motor system. Unlike conventional spherical robots, MonoRollBot employs a minimalist actuation approach, relying on only one motor and a passive spring to control its locomotion. The robot achieves 3-DOF motion through an innovative coupling of spring dynamics and motor control. In this work, we detail the design of the MonoRollBot and evaluate its motion capabilities through design studies. We also do studies on its locomotion behaviours based on changes in rotating mass and stiffness properties.
@@ II-A Overall Structure
The robot, referred to as MonoRollBot, is a 3-DOF spherical robot designed to achieve rolling and balancing within a spherical shell as shown in Fig. 1. The robot’s body integrates both a linear actuator and a rotating mass system, forming the core mechanism for movement and stability on $\Sigma_{l}$ frame. The key design goal is to enable spherical motion with an underactuated mechanism, utilizing fewer motors (only one motor here) to achieve efficient movement of mass-point on $\Sigma_{m}$ frame. The robot consists of a rolling shell with the frame of $\Sigma_{s}$ and an internal driving structure. The external rolling shell encapsulates the internal components, providing protection and omnidirectional motion on a flat surface. The movement is characterized by a global coordinate system defined by $\Sigma_{o}$.
The aim is to rotate a symmetric mass in multiple directions to create 3DoF capabilities for MonoRollBot. The robot uses a combination of a linear actuator and a rotating spring-mass system to generate movement. Table I presents all the mechanical parts and their corresponding parametric variables. This design reduces the number of active motors, achieving actuation through a single spring-motor combination. For the linear actuator body, the linear actuator is arranged along the central vertical axis of the robot, driving vertical motion and assisting in internal mass transfer as shown in Fig. 2. The positioning of the actuator allows the centre of mass to be changed, thus affecting the rolling direction of the spherical shell. In addition, the rotating spring mass is connected to a spring-loaded mechanism that stores potential energy and releases it to assist in motion. The mass rotates around the central axis of the robot, affecting its orientation and balance, and is therefore essential for generating rotational motion.
@@ II-B Robot Kinematics and Driving Mechanism Model
where $C$ and $S$ are the cosine functions. Next, to determine the amount of locomotion that happens by the rolling sphere on the plane $(x,y)$, we used linear kinematics to determine the sphere contact location
@@
Next, we need to estimate the location of the rotating mass within the sphere, as direct sensory feedback is unavailable and the rotation speed is too fast for inertial sensors to track accurately. To address this, we utilize motor encoder data and motion dynamics to determine the position of the rotating spiral nut gear. Initially, the motor torque is directly related to the springs connected to the spherical mass points including other connecting materials $m_{b}=m+m_{sb}$, with the following motion equation formulated for discrete-time computation [12] by
@@
${\bm{\mathbf{T}}}_{d}={\bm{\mathbf{R}}}\left[\begin{array}[]{c}0\\ 0\\ d_{d}(k)\end{array}\right],{\bm{\mathbf{T}}}_{s}={\bm{\mathbf{R}}}\left[\begin{array}[]{c}R\sin\theta_{d}\;\cos\theta_{n}\\ R\sin\theta_{d}\;\sin\theta_{n}\\ d_{d}(k)\end{array}\right],$ (5)
where $\theta_{d}=2\sin^{-1}(d_{d}/2R)$. Then, the angle $\alpha$ can be calculated by the dot product of vectors in (5) as
@@ III Motion Study
This section examines the capabilities of the designed robot from two perspectives: variations in rotating mass and changes in the stiffness of the connected mass. We also analyze factors such as overall behavior, total generated force, motor response, and additional parameters to assess the potential of MonoRollBot.
@@ III-A Rotating Mass Diversity
To investigate the effect of different rotating masses, we conducted experiments with a series of the mass of the internal rotating component varied between $\left[20-70\right]$ g. The robot was placed on a flat surface with the same position for every experiment, and the internal motor was actuated to induce spherical motion with pre-defined length of motion on $d_{a}$ (being from initial to final length of spiral lead). Throughout the experiment, an IMU sensor fixed inside the robot measured the rotational angles in the global coordinate system $\left(\gamma,\beta,\alpha\right)$, while the motor encoder recorded the motor’s angular rotation $\left(\theta_{m},\dot{\theta}_{m}\right)$
@@
Fig. 4. shows the robot’s rotational angles through IMU and its path trajectory ${\bm{\mathbf{D}}}$ under different mass conditions based on Eq. (2). For lighter masses ($20$, $35$) g, the robot’s rotational response remains relatively linear motion, with smoother and less oscillatory angles $\left(\gamma,\beta,\alpha\right)$. As seen in the trajectory plots, the robot follows a more predictable and controlled path, as the lighter mass requires less effort to maintain balance and motion. The lower inertia allows the robot to adjust its orientation without causing excessive tilting or abrupt changes in direction but to propel it requires a certain conservation of angular momentum. As the mass increases ($50,70$) g, the rotational angles show more pronounced oscillations, especially in the $\gamma$-axis. This suggests that the larger mass creates more dynamic dependency, making the robot’s motion requireing more complex dynamic based controllers especially when mass moves upper hemisphere as shown in Fig. 5 (estimated by Eq. (4)). In the trajectory plots, heavier masses result in larger deviations from a straight-line path, with more erratic turns and fluctuations. The added inertia from the larger mass amplifies the robot’s response to motor actuation, leading to quicker changes in direction and less smooth trajectories.

### v3

@@ Abstract
Spherical rolling robots have garnered significant attention in the field of mobile robotics for applications such as inspection and space exploration. Designing underactuated rolling robots poses challenges in achieving multi-directional propulsion with high degrees of freedom while utilizing a limited number of actuators. This paper presents the MonoRollBot, a novel 3-degree-of-freedom (DOF) spherical robot that utilizes an underactuated mechanism driven by only a single spring-motor system. Unlike conventional spherical robots, MonoRollBot employs a minimalist actuation approach, relying on only one motor and a passive spring to control its locomotion. The robot achieves 3-DOF motion through an innovative coupling of spring dynamics and motor control. In this work, we detail the design of the MonoRollBot and evaluate its motion capabilities through design studies. We also study its locomotion behaviours based on changes in rotating mass and stiffness properties.
@@ II-A Overall Structure
The robot, referred to as MonoRollBot, is a 3-DOF spherical robot designed to achieve rolling and balancing within a spherical shell as shown in Fig. 1. All components of the robot, except for the motor, are fabricated using 3D printing technology. Polylactic acid (PLA) is used as the printing material due to its mechanical efficiency and sufficient strength, which are essential for maintaining the structural integrity and operation of the robot while minimizing its weight. The robot’s body integrates both a linear actuator and a rotating mass system, forming the core mechanism for movement and stability on $\Sigma_{l}$ frame. The key design goal is to enable spherical motion with an underactuated mechanism, utilizing fewer motors (only one motor here) to achieve efficient movement of mass-point on $\Sigma_{m}$ frame. The robot consists of a rolling shell with the frame of $\Sigma_{s}$ and an internal driving structure. The external rolling shell encapsulates the internal components, providing protection and omnidirectional motion on a flat surface. The movement is characterized by a global coordinate system defined by $\Sigma_{o}$.
The aim is to rotate a symmetric mass in multiple directions to create 3DoF capabilities for MonoRollBot. The robot uses a combination of a linear actuator and a rotating spring-mass system to generate movement. Table I presents all the mechanical parts and their corresponding parametric variables. The parameters in Table I were determined through theoretical analysis, experimental validation, and design constraints considering single DC motor. Key considerations include optimizing the spiral screw pitch $P$ for balanced torque $\tau_{m}$ and linear displacement and selecting the motor’s nominal current to meet load and power requirements. These values were iteratively refined via simulations and prototype testing to ensure optimal performance for moving mass along the sliding bar. This design reduces the number of active motors, achieving actuation through a single spring-motor combination. The motor, positioned at the central axis of the robot, drives the rotating nut through a gear system. This nut, in turn, interacts with the spiral lead to generate both linear and rotational motion as shown in Fig. 2. The spherical mass, mounted on the rotating nut, does not spin at the same speed as the spiral lead; instead, its motion by $\theta_{n}$ angle is influenced by the spring dynamics and the resultant motor torque through gears. For the linear actuator body, the linear actuator is arranged along the central vertical axis of the robot, driving vertical motion and assisting in internal mass transfer as shown in Fig. 2. The positioning of the actuator allows the centre of mass to be changed, thus affecting the rolling direction of the spherical shell. In addition, the rotating spring mass is connected to a spring-loaded mechanism that stores potential energy and releases it to assist in motion. The mass rotates around the central axis of the robot, affecting its orientation and balance, and is therefore essential for generating rotational motion.
@@
Figure 3: Actuator operation captured in snapshots where the mass rotates from the bottom of the screw lead to the midpoint of the sphere.
@@ II-B Robot Kinematics and Driving Mechanism Model
where $C$ and $S$ are the cosine functions. Next, to obtain the amount of locomotion that happens by the rolling sphere on the plane $(x,y)$, we used linear kinematics to calculate the sphere contact location
@@
Next, we need to estimate the location of the rotating mass within the sphere, as direct sensory feedback is unavailable and the rotation speed is too fast for inertial sensors to track accurately. To address this, we utilize motor encoder data and motion dynamics to determine the position of the rotating spiral nut gear. Initially, the motor torque is directly related to the springs connected to the spherical mass points including other connecting materials $m_{b}=m+m_{sb}$, with the following motion equation formulated for discrete computation [12] by
@@
${\bm{\mathbf{T}}}_{d}={\bm{\mathbf{R}}}\left[\begin{array}[]{c}0\\ 0\\ d_{a}(k)\end{array}\right],{\bm{\mathbf{T}}}_{s}={\bm{\mathbf{R}}}\left[\begin{array}[]{c}R\sin\theta_{d}\;\cos\theta_{n}\\ R\sin\theta_{d}\;\sin\theta_{n}\\ d_{a}(k)\end{array}\right],$ (5)
where $\theta_{d}=2\sin^{-1}(d_{a}/2R)$. Then, the angle $\alpha$ can be calculated by the dot product of vectors in (5) as
@@ III Motion Study
This section examines the capabilities of the designed robot from two perspectives: variations in rotating mass and changes in the stiffness of the connected mass. We also analyze factors such as overall behavior, total generated force, motor response, and additional parameters to assess the potential of MonoRollBot.The experiments in this section use predefined input signals to test the robot’s ability to maintain a stable trajectory while experiencing sudden changes in angular velocity and torque. The data collected in the experiments include motion trajectories, angular displacements, and system responses, as shown in Fig. 4.
@@ III-A Rotating Mass Diversity
To investigate the effect of different rotating masses, we conducted experiments with a series of the mass of the internal rotating component varied between $\left[20-70\right]$ gram. The robot was placed on a flat surface with the same position for every experiment, and the internal motor was actuated to induce spherical motion with pre-defined length of motion on $d_{a}$ (being from initial to final length of spiral lead). Throughout the experiment, an IMU sensor fixed inside the robot measured the rotational angles in the global coordinate system $\left(\gamma,\beta,\alpha\right)$, while the motor encoder recorded the motor’s angular rotation $\left(\theta_{m},\dot{\theta}_{m}\right)$
The motor speed was set to a fixed value during all experiments. The control process involved sending predefined speed commands to motor drivers through the communication interface. This approach ensured consistent and repeatable actuation during the experiments, eliminating variations caused by the dynamic control algorithm. During the operation of the experiment, the motors drove the helical lead mechanism at a specified speed. This simple control method allowed for a controlled study of the robot’s kinematic dynamics under different mass and stiffness configurations, isolating the effects of the physical parameters from the complex control effects.
@@
Fig. 4. shows the robot’s rotational angles through IMU and its path trajectory ${\bm{\mathbf{D}}}$ under different mass conditions based on Eq. (2). For lighter masses ($20$, $35$) g, the robot’s rotational response remains relatively linear motion, with smoother and less oscillatory angles $\left(\gamma,\beta,\alpha\right)$. As seen in the trajectory plots, the robot follows a more predictable and controlled path, as the lighter mass requires less effort to maintain balance and motion. The lower inertia allows the robot to adjust its orientation without causing excessive tilting or abrupt changes in direction but to propel it requires a certain conservation of angular momentum. As the mass increases ($50,70$) g, the rotational angles show more pronounced oscillations, especially in the $\gamma$-axis. This suggests that the larger mass creates more dynamic dependency, making the robot’s motion requiring more complex dynamic-based controllers especially when mass moves the upper hemisphere as shown in Fig. 5 (estimated by Eq. (4)). In the trajectory plots, heavier masses result in larger deviations from a straight-line path, with more erratic turns/swings and fluctuations. The added inertia from the larger mass amplifies the robot’s response to motor actuation, leading to quicker changes in direction or backward motion (which is minor-based on expected direction) and less smooth trajectories. In Experiment 3, the internal mass rotated at a constant velocity (Fig. 6), while the robot’s actuator drove the system. Key snapshots highlight the interaction between the spherical mass, actuator, and spiral lead mechanism. Observed trajectory deviations and angular shifts reveal the impact of mass distribution on the robot’s motion dynamics and control performance.